When the radius of a capacitor is halfed what will be its resistance?
1
Expert's answer
2020-06-22T11:10:12-0400
"R'=\rho \frac{l}{\pi (0.5r)^2}=4\rho \frac{l}{\pi r^2}"

"R'=4R"